# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ 6 Pseudocodes
Algorithm 1 Dispersion in $1$-TVG with $n+1$ agents
1 if $a_{.}state=explore$ then
2       call Algorithm 4
3else
4       call Algorithm 5
Algorithm 2 Movement of agent $a_{i}\in G_{2}$ at $v$ with $a_{i}.settled=0$
@@
14                                     set $a_{i}.state=backtrack$
15                                     move through $a_{i}.prt\_out$
@@
20                         if $label=a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
21                               $a_{i}.state=backtrack$
22                               move through $a_{i}.prt\_in$
23                        else if $label\neq a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
@@
27                         set $a_{i}.divide=1$
28                         call $Group\_divide()$ (Algorithm 6)
29
@@
33                               call $Group\_divide()$ (Algorithm 6)
34
@@
38             if the movement of agent $a_{i}$ in round $r-1$ is successful then
@@
42else if $a_{i}.state=backtrack$ then
@@
46                   if $parent$ component of $r_{j}^{v}(G_{1})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
@@
48                         set $a_{i}.state=explore$
49                         move through $a_{i}.prt\_out$
50                  else if $parent$ component of $r_{j}^{v}(G_{1})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
51
@@
53                         if $a_{i}.prt\_out=parent$ and $parent$ is the component of $r_{j}^{v}(G_{1})$, where $r_{j}$ is the settled agent at node $v$ then
@@
56                               set $a_{i}.state=explore$
57                               move through $a_{i}.prt\_out$
@@
60                         set $a_{i}.divide=1$
61                         call $Group\_divide()$ (Algorithm 6)
62
@@
66                               call $Group\_divide()$ (Algorithm 6)
67
@@
71             if the movement of agent $a_{i}$ in round $r-1$ is successful then
@@
16                   if $a_{i}.grp\_label=label$ and $a_{i}.dfs\_label=dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
@@
20                  else if $a_{i}.grp\_label\neq label$ or $a_{i}.dfs\_label\neq dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
@@
29                   if $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
@@
31                               $a_{i}.count_{2}=0$
32                               $a_{i}.skip=-1$
33                               $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
34                               $a_{i}.prt\_out=0$
@@
40                  else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
@@
44                               $a_{i}.state=backtrack$
45                               move through $a_{i}.prt\_out$
@@
55                               set $a_{i}.state=explore$
56                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
57                               set $a_{i}.skip=a_{i}.prt\_out$
@@
61                               call $Group\_divide()$ (Algorithm 6)
62
63else if $r$ mod 2=0 then
64       if the movement of agent $a_{i}$ in round $r-1$ is successful then
@@
1 if $r$ mod 2 =1 then
2       if $a_{i}.success=True$ then
3             set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
4             set $a_{i}.count_{2}=0$
5             if  $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
6                   if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
7                         $a_{i}.state=explore$
8                         $a_{i}.skip=-1$
9                         $a_{i}.prt\_out=0$
10                         $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
11                         move through $a_{i}.prt\_out$
13                         if $a_{i}.prt\_out=a_{i}.skip$ then
14                               $a_{i}.prt\_out=(a_{i}.prt\_out+1)\mod\delta_{v}$
15                               if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
16                                     $a_{i}.state=explore$
17                                     $a_{i}.skip=-1$
18                                     $a_{i}.prt\_out=0$
19                                     $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
20                                     move through $a_{i}.prt\_out$
22                                     $a_{i}.state=explore$
23                                     move through $a_{i}.prt\_out$
24                        else if $a_{i}.prt\_out\neq a_{i}.skip$ then
25                               $a_{i}.state=explore$
26                               move through $a_{i}.prt\_out$
27            else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
28                   if $a_{i}.prt\_out=parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
29                         move through $a_{i}.prt\_out$
30                  else if $a_{i}.prt\_out\neq parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
31                         $a_{i}.state=explore$
32                         move through $a_{i}.prt\_out$
33      else if $a_{i}.success=False$ then
34
36                   $a_{i}.count_{2}=0$
37                   set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
38                   set $a_{i}.state=explore$
39                   $a_{i}.skip=a_{i}.prt\_out$
41                   move through $a_{i}.prt\_out$
43                   $a_{i}.count_{2}=a_{i}.count_{2}+1$
44                   if $a_{i}.count_{2}=4n^{2}$ then
45                         set $a_{i}.count_{2}=0$, and $a_{i}.count_{3}=a_{i}.count_{3}+1$
46                         if $a_{i}.count_{3}<4n^{2}$ then
47                               set $a_{i}.state=explore$
48                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
49                               set $a_{i}.skip=a_{i}.prt\_out$
51                               move through $a_{i}.prt\_out$
52                        else
53                               call $Group\_divide()$ (Algorithm 6)
54
55                  else if $a_{i}.count_{2}<4n^{2}$ then
56                         $a_{i}.count_{2}=a_{i}.count_{2}+1$
57                         move through $a_{i}.prt\_out$
58else if $r$ mod 2 = 0 then
59       if the movement of agent $a_{i}$ in round $r-1$ is successful then
60             set $a_{i}.success=True$
61      else
62             set $a_{i}.success=False$
Algorithm 5 $a_{i}.state=backtrack$
2 if $i\leq\lceil{\frac{x}{2}}\rceil$ then
[…]
4       $a_{i}.prt\_out=0$
5       set $a_{i}.state=explore$
6       $a_{i}.skip=-1$
7       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
8       move through $a_{i}.prt\_out$
9else
[…]
11       $a_{i}.prt\_out=0$ set $a_{i}.state=explore$
12       $a_{i}.skip=-1$
13       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
14       move through $a_{i}.prt\_out$
Algorithm 6 $Group\_divide()$
@@
6                   if $label$ component of $a_{i}^{v}(G_{1})$ does not match with $a_{j}.grp\_label$ then
@@
8                  else if $label$ component of $a_{i}^{v}(G_{1})$ matches with $a_{j}.grp\_label$ then

### v3

@@ 8.1 Dispersion with $n+1$ co-located agents
Algorithm 1 Dispersion in 11-TVG with $n+1$ agents
8.1.1 The algorithm
In this section, we provide a detailed description of the algorithm for dispersion in 1-bounded 1-interval connected graph $\mathcal{G}$. In our algorithm, unsettled agents may get divided into two groups, which we call $G_{1}$ and $G_{2}$ respectively. At any particular round, there can be no more than two groups. However, over time, if all the unsettled agents of one group get settled, the other group can be divided into two groups again. So, if we look throughout the run of the algorithm, there may exist several groups. To identify these groups, we provide unique labelling to each group whenever two groups are formed, and they store their new ID in a parameter. Now we provide the parameters maintained by each agent $a_{i}$.
@@ 8.1.1 The algorithm
14                  set $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$
@@
19          if $label=a_{i}.grp\_label$, where labellabel is the component of $r_{j}^{v}(G_{1})$ then
20             $a_{i}.state=backtrack$ and move through $a_{i}.prt\_in$
21          else if $label\neq a_{i}.grp\_label$, where labellabel is the component of $r_{j}^{v}(G_{1})$ then
@@
25            set $a_{i}.divide=1$ and call Group_divide()Group\_divide()
@@
29               call Group_divide()Group\_divide()
@@
33    if the movement of agent $a_{i}$ in round r1r-1 is successful then
@@
37else if $a_{i}.state=backtrack$ then
@@
41       if parentparent component of $r_{j}^{v}(G_{1})$ is 1-1, where $r_{j}$ is the settled agent at node $v$ then
@@
43            set $a_{i}.state=explore$ and move through $a_{i}.prt\_out$
44       else if parentparent component of $r_{j}^{v}(G_{1})$ is not 1-1, where $r_{j}$ is the settled agent at node $v$ then
45
@@
47          if $a_{i}.prt\_out=parent$ and parentparent is the component of $r_{j}^{v}(G_{1})$, where $r_{j}$ is the settled agent at node $v$ then
@@
50               set $a_{i}.state=explore$ and move through $a_{i}.prt\_out$
@@
53            set $a_{i}.divide=1$ and call Group_divide()Group\_divide()
@@
57               call Group_divide()Group\_divide()
@@
61    if the movement of agent $a_{i}$ in round r1r-1 is successful then
@@
2 if $i\leq\lceil{\frac{x}{2}}\rceil$ then
3   update $a_{i}.grp\_label=a_{i}.grp\_label.0$
4 $a_{i}.prt\_out=0$
5   set $a_{i}.state=explore$
6 $a_{i}.skip=-1$
7   set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
8   move through $a_{i}.prt\_out$
9else
10   update $a_{i}.grp\_label=a_{i}.grp\_label.1$
11 $a_{i}.prt\_out=0$ set $a_{i}.state=explore$
12 $a_{i}.skip=-1$
13   set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
14   move through $a_{i}.prt\_out$
Algorithm 3 Group_divide()Group\_divide()
  • 1.
  • 2.
  • 3.
  • 4.
  • 5.
  • 6.
    $\bm{a_{i}.success}$: When an agent $a_{i}$ attempts to move from a node $u$ to $v$ via the edge (u,v)(u,\,v), in the round $t$, it is either successful or unsuccessful. If the agent reaches the node $v$ in the round $t+1$, it is a successful attempt, thus, $a_{i}.success=True$; otherwise, $a_{i}.success=False$. Each agent $a_{i}$ initially has $a_{i}.success=True$.
  • 7.
    $\bm{a_{i}.skip}$: If agent $a_{i}$ skips a port, then it stores it in this parameter. Initially, $a_{i}.skip=-1$. This parameter allows agents of $G_{2}$ to store port numbers when they skip at the root of their DFS algorithm.
  • 8.
  • 9.
    $\bm{a_{i}.grp\_label}$: It stores a binary string. Initially $a_{i}.grp\_label=10$. If $a_{i}.grp\_label$ ends with 0, it implies $a_{i}$ is a part of $G_{1}$. Else if $a_{i}.grp\_label$ ends with 1, it implies $a_{i}$ is a part of $G_{2}$. Later, we elaborate on how this variable gets updated.
  • 10.
    $\bm{a_{i}.count_{1}}$: If agent $a_{i}\in G_{1}$, then it stores the number of consecutive odd rounds with $a_{i}.success=False$. Initially, each agent has $a_{i}.count_{1}=0$. This parameter is used by agents in $G_{1}$ to determine when $G_{2}$ has been fully dispersed and when $G_{1}$ can be further divided into two new groups.
  • 11.
    $\bm{a_{i}.count_{2}}$: If agent $a_{i}\in G_{2}$ then it stores the number of consecutive odd rounds with $a_{i}.success=False$. This parameter is used by agents in $G_{2}$ to determine when $G_{2}$ has needs
  • 12.
    $\bm{a_{i}.count_{3}}$: If agent $a_{i}\in G_{2}$, then it stores the number of new DFS traversals it started from the time $a_{i}$ became a part of this group. This parameter is used by agents in $G_{2}$ to determine when $G_{1}$ has been fully dispersed and when $G_{2}$ can be further divided into two new groups.
Let $a_{i}$ be a settled agent at node $v$. Any unsettled agent belongs to either $G_{1}$ or $G_{2}$. When an unsettled agent $a_{j}$ while doing its movement (along with other unsettled agents, if any, in its group) according to our algorithm, reaches a node $v$ where an agent $a_{i}$ is already settled, $a_{j}$ needs some information regarding $v$ from $a_{i}$ to continue its traversal. In this regard, each settled agent maintains two parameters, one with respect to each of the groups as mentioned below.
  • 1.
    $\bm{a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)}$: Initially $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(-1,\,-1,\,0)$. Let $t$ be the round when agent $a_{i}$ settles at node $v$. Let $t^{\prime}\geq t$ be the last round when $a_{i}$ saw some agent of $G_{1}$, say, agent $a_{j}$ at node $v$. In $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)$, $a_{i}$ stores $(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$. That is, it stores the information regarding node $v$ w.r.t. the DFS traversal of $G_{1}$ in round $t^{\prime}$. Note that throughout this paper, we denote $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)$ by $a_{i}^{v}(G_{1})$.
  • 2.
    $\bm{a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)}$: Initially $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,-1,\,0)$. Let $t$ be the round when agent $a_{i}$ settles at node $v$. Let $t^{\prime}\geq t$ be the last round when $a_{i}$ saw some agent of $G_{2}$, say, agent $a_{j}$ at node $v$. In $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)$, $a_{i}$ stores $(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$. That is, it stores the information regarding node $v$ w.r.t. the DFS traversal of $G_{2}$ in round $t^{\prime}$. Note that throughout this paper, we denote $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)$ by $a_{i}^{v}(G_{2})$.
Suppose $n+1$ agents are co-located at a node, say $v$, of the graph (rooted configuration). Initially, for each agent $a_{i}$, $a_{i}.grp\_label=10$. Therefore, all agents are a part of $G_{1}$. According to our model, an agent cannot sense the presence of a missing edge unless it attempts to move through it. To tackle this, we take two consecutive rounds to perform one edge traversal. In every odd round, agents move, and in the next even round, it realizes the movement is successful or unsuccessful. Precisely, say an agent $a_{i}$ attempts to move in an odd round $r$. If its movement is successful, it updates its parameter $a_{i}.success$ to $True$ in the even round $r+1$. Otherwise, it updates $a_{i}.success=False$ in the even round $r+1$. Now we proceed with a detailed description of our algorithm (Refer to Algorithm 1 for the main algorithm).
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{1}}$ at odd round: In any odd round, let agent $a_{i}\in G_{1}$ be at node $u$ and $a_{i}.settled=0$. Based on $a_{i}.divide$, it follows the following steps.
  • 1.
    $\bm{a_{i}.divide=0}$: Suppose, agent $a_{i}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ (lines 5-8 of Algorithm 2). Further, it updates $a_{i}^{w}(G_{1})$. Else if agent $a_{i}$ is not the minimum ID, then it moves according to the DFS traversal algorithm (lines 11-17 of Algorithm 2). If agent $a_{i}$ reaches a node $v$ and finds an agent with $a_{j}$, with $a_{j}.settled=1$, then $a_{i}$ does not settle at node $v$ and moves according to the DFS traversal algorithm (lines 18-24 and lines 42-54 of Algorithm 2). In this way, the agent $a_{i}\in G_{1}$ continues the DFS traversal of the graph unless $a_{i}$ encounters a missing edge for the first time. In particular, at some odd round $r$, agent $a_{i}$ moves via an edge $e$, and the movement is not successful at the end of round $r$. This is understood by agent $a_{i}$ in the round $r+1$. In round $r+2$, agent $a_{i}$ updates dividedivide by 1 (lines 23-25 and lines 50-52 of Algorithm 2) and divides into two groups using the Group_divideGroup\_divide procedure as follows.
    $\bm{Group\_divide}$: (refer to Algorithm 3) Half of the agents with smaller IDs at node $u$ are part of $G_{1}$ and the other half of the agents at node $u$ are part of $G_{2}$. If agent $a_{i}$ is part of $G_{2}$, then sets $a_{i}.grp\_label=a_{i}.grp\_label.1$ by appending 1 at the end and starts a new DFS traversal considering the current node as the new root. In this case, agent $a_{i}$ also updates the other parameters $a_{i}.skip=-1$, $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$ and $a_{i}.prt\_out=0$. Else if agent $a_{i}$ remains as a part of $G_{1}$, then sets $a_{i}.grp\_label=a_{i}.grp\_label.0$ by appending 0 at the end and starts a new DFS traversal considering the current node as the new root. In this case, agent $a_{i}$ also updates the other parameters $a_{i}.skip=-1$, $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$ and $a_{i}.prt\_out=0$. Agent $a_{i}$ moves through $a_{i}.prt\_out$.
  • 2.
    $\bm{a_{i}.divide=1}$: (refer to Algorithm 2) Suppose, agent $a_{i}\in G_{1}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{2}$ at node $w$, then $a_{i}$ updates $a_{i}^{u}(G_{2})$ to $(a_{j}.prt\_in,a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}\in G_{1}$ reaches a node $w$ and find agent $a_{j}$ with $a_{j}.settled=1$, then it does not settle at the node $v$ and it can understand whether node $w$ is the new node according to the current DFS or previously visited node by comparing labellabel, dfs_labeldfs\_label components of $a_{j}^{w}(G_{1})$ with $a_{i}.grp\_label$, $a_{i}.dfs\_label$ respectively. If both matches then node $w$ is visited previously in the current DFS traversal of agent $a_{i}$. Otherwise, if any of those two components do not match, then agent $a_{i}\in G_{1}$ treats $w$ as an unvisited node w.r.t. the current DFS. Round $r+2$ onwards, $a_{i}$ updates its $count_{1}$ parameter as follows. At odd round $r_{1}\geq r+2$, if it finds $a_{i}.success=True$, i.e. it realized a successful movement in the even round $r_{1}-1$, agent $a_{i}$ updates $a_{i}.count_{1}=0$; else if $a_{i}.success=False$, agent $a_{i}$ updates $a_{i}.count_{1}=a_{i}.count_{1}+1$. If at some odd round, $a_{i}$ finds $a_{i}.count_{1}=16n^{2}$, then unsettled agents at node $u$ divide into two groups again using aforementioned Group_divideGroup\_divide procedure.vivivi Note that all unsettled agents at node $u$ belong are from $G_{1}$. We will show in the analysis that once the $count_{1}$ parameter for agents in $G_{1}$ reaches $16n^{2}$, all agents in $G_{2}$ becomes settled agent. That is, $G_{1}$ divides into two groups only after getting stuck at a node for a particular port for $16n^{2}$ many consecutive odd rounds.
1 if $r$ mod 2 =1 then
2 if $a_{i}.success=True$ then
3      set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$ and $a_{i}.count_{2}=0$
4    if  parentparent component of $r_{j}^{v}(G_{2})$ is 1-1, where $r_{j}$ is the settled agent at node $v$ then
5       if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
6          $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, and $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
7            move through $a_{i}.prt\_out$
9          if $a_{i}.prt\_out=a_{i}.skip$ then
10             $a_{i}.prt\_out=(a_{i}.prt\_out+1)\mod\delta_{v}$
11             if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
12                $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, and $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
13                  move through $a_{i}.prt\_out$
15                $a_{i}.state=explore$ and move through $a_{i}.prt\_out$
16          else if $a_{i}.prt\_out\neq a_{i}.skip$ then
17             $a_{i}.state=explore$ and move through $a_{i}.prt\_out$
18    else if parentparent component of $r_{j}^{v}(G_{2})$ is not 1-1, where $r_{j}$ is the settled agent at node $v$ then
19       if $a_{i}.prt\_out=parent$, where parentparent is the component of $r_{j}^{v}(G_{2})$ then
20            move through $a_{i}.prt\_out$
21       else if $a_{i}.prt\_out\neq parent$, where parentparent is the component of $r_{j}^{v}(G_{2})$ then
22          $a_{i}.state=explore$ and move through $a_{i}.prt\_out$
23 else if $a_{i}.success=False$ then
24
26       $a_{i}.count_{2}=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.state=explore$, and $a_{i}.skip=a_{i}.prt\_out$
28         move through $a_{i}.prt\_out$
30       $a_{i}.count_{2}=a_{i}.count_{2}+1$
31       if $a_{i}.count_{2}=4n^{2}$ then
32            set $a_{i}.count_{2}=0$, and $a_{i}.count_{3}=a_{i}.count_{3}+1$
33          if $a_{i}.count_{3}<4n^{2}$ then
34               set $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and $a_{i}.skip=a_{i}.prt\_out$
36               move through $a_{i}.prt\_out$
37          else
38               call Group_divide()Group\_divide()
39       else if $a_{i}.count_{2}<4n^{2}$ then
40          $a_{i}.count_{2}=a_{i}.count_{2}+1$
41            move through $a_{i}.prt\_out$
42else if $r$ mod 2 = 0 then
43 if the movement of agent $a_{i}$ in round r1r-1 is successful then
44      set $a_{i}.success=True$
45 else
46      set $a_{i}.success=False$
Algorithm 4 $a_{i}.state=backtrack$
@@
16       if $a_{i}.grp\_label=label$ and $a_{i}.dfs\_label=dfs\_label$, where labellabel and dfs_labeldfs\_label are the component from $r_{j}^{v}(G_{2})$ then
@@
20       else if $a_{i}.grp\_label\neq label$ or $a_{i}.dfs\_label\neq dfs\_label$, where labellabel and dfs_labeldfs\_label are the component from $r_{j}^{v}(G_{2})$ then
@@
29       if parentparent component of $r_{j}^{v}(G_{2})$ is 1-1, where $r_{j}$ is the settled agent at node $v$ then
@@
31             $a_{i}.count_{2}=0$, $a_{i}.skip=-1$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=0$
@@
36       else if parentparent component of $r_{j}^{v}(G_{2})$ is not 1-1, where $r_{j}$ is the settled agent at node $v$ then
@@
40             $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$
@@
50               set $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$
@@
54               call Group_divide()Group\_divide()
55else if $r$ mod 2=0 then
56 if the movement of agent $a_{i}$ in round r1r-1 is successful then
@@
1 if $a_{.}state=backtrack$ then
2   call Algorithm 4
3else
4   call Algorithm 5
Algorithm 6 Movement of agent $a_{i}\in G_{2}$ at $v$ with $a_{i}.settled=0$
  • (1)
    If $\bm{a_{i}.state=explore}$ and $\bm{a_{i}.success=True}$: It updates $a_{i}.count_{2}=0$ and does the following (lines 2-26 of Algorithm 5).
    • (a)
      Node $\bm{u}$ has no agent with $\bm{settled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. Note that in this case, there is no agent from $G_{1}$, since in the Group_divideGroup\_divide procedure, the first half of the agents with smaller IDs are assigned to $G_{1}$.
      If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
    • (b)
      Node $\bm{u}$ has an agent with $\bm{settled=1}$: Let $a_{j}$ be at node $u$ with $a_{j}.settled=1$. If $a_{i}.grp\_label$, $a_{i}.dfs\_label$ matches labellabel, dfs_labeldfs\_label components of $a_{j}^{u}(G_{2})$ respectively, then it is a previously visited node. In this case, $a_{i}$ sets $a_{i}.state=backtrack$, $a_{i}.prt\_out=a_{i}.prt\_in$, and it moves through $a_{i}.prt\_out$. Otherwise, it is an unvisited node w.r.t. the current DFS of $a_{i}$. In this case, it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
  • (2)
    If $\bm{a_{i}.state=explore}$ and $\bm{a_{i}.success=False}$: It does the following (lines 27-54 of Algorithm 5).
    • (a)
      • (i)
        ($\alpha$) $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If $a_{i}.prt\_out=\delta_{u}-1$, then agent $a_{i}$ sets $a_{i}.state=explore$, stores $a_{i}.skip=-1$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$ (i.e., start a new DFS traversal), $a_{i}.prt\_out=0$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq\delta_{u}-1$, then it sets $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{u}$ and sets $a_{i}.count_{2}=0$ and move through $a_{i}.prt\_out$.
      • (ii)
    • (b)
      There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using Group_divideGroup\_divide procedure mentioned in the algorithm for unsettled agents of $G_{1}$. Note that if $a_{i}.count_{3}=4n^{2}$, then $G_{1}$ is already dispersed as justified in the correctness.
  • (3)
    If $\bm{a_{i}.state=backtrack}$ and $\bm{a_{i}.success=True}:$ It updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{u}$ and $a_{i}.count_{2}=0$. Let $a_{j}$ be the agent at node $u$ with $a_{j}.settled=1$. Based on parentparent component of $a_{j}^{u}(G_{2})$, there are two cases (lines 2-22 of Algorithm 4).
    • (a)
      $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If agent $a_{i}$ finds parentparent component of $a_{j}^{u}(G_{2})$ to be 1-1, then node $u$ is the root node of the current DFS traversal of agent $a_{i}$. In this case, if $a_{i}.prt\_out=$ minimum available port except for $a_{i}.skip$, then it sets $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq$ minimum available port except for $a_{i}.skip$, then it does the following.
    • (b)
      $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: If agent $a_{i}$ does not find parentparent component of $a_{j}^{u}(G_{2})$ to be 1-1, then node $u$ is not the root node of the current DFS traversal of agent $a_{i}$. If $a_{i}.prt\_out$ matches with parentparent component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out$ does not match with parentparent component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
@@
    6       if labellabel component of $a_{i}^{v}(G_{1})$ does not match with $a_{j}.grp\_label$ then
@@
    8       else if labellabel component of $a_{i}^{v}(G_{1})$ matches with $a_{j}.grp\_label$ then